OpenStudy (inkyvoyd):

Just a note to myself \(\Huge \lim_{x\rightarrow 1.8835}\frac{f(x+\delta x-f(x)}{\delta x}\)

OpenStudy (inkyvoyd):

\(\Huge \lim_{x\rightarrow 1.8835}\frac{f(x+\Delta x)-f(x)}{\Delta x}\)

OpenStudy (inkyvoyd):

\(\Huge \lim_{t\rightarrow 1.8835}\frac{56.76-16(t+\Delta t)^2-56.76+16t^2}{\Delta t}\)

OpenStudy (inkyvoyd):

\(\Huge \lim_{t\rightarrow 1.8835}\frac{f(t+\Delta t)-f(t)}{\Delta t}\)

OpenStudy (inkyvoyd):

\(\Huge \lim_{\Delta t\rightarrow 0}\frac{-32t\Delta t-16\Delta t^2}{\Delta t}\)

OpenStudy (inkyvoyd):

\(\Huge \lim_{\Delta t\rightarrow 0}\frac{f(t+\Delta t)-f(t)}{\Delta t}\)